OpenStudy (anonymous):

I'm trying to verify this expression: 1 + sec^2 x sin^2 x = sec^2 x This is what I've done: sec^2 x =1+tan^2 x =sin^2 x +cos^2 x +tan^2 x =(sin^2 x cosx/cosx)+(cos^2x sinx/sinx)+(sinx/cosx) =sinx[sinx cosx/cosx)+(cos^2 x/sinx)+(1/cosx) I'm pretty sure I'm doing this all wrong! Please help!

OpenStudy (anonymous):

i made a mistake before so now is the correct one secx=1/cosx and sec^2xsin^2x=tan^2x so 1+tan^2x=sec^2x

OpenStudy (anonymous):

Sorry I left for a few minutes I have to match the right side to the left side

OpenStudy (sirm3d):

how about changing \(\sec^2 x\) to \(\displaystyle \frac{1}{\cos^2 x}\)?

OpenStudy (sirm3d):

@mikoza already gave the solution.

OpenStudy (anonymous):

Isn't that matching the left side to the right?

OpenStudy (sirm3d):

yes it is.

OpenStudy (anonymous):

I'm supposed to change the right side to match the left. I know the other one is easier, but I'm having trouble matching the right to left

OpenStudy (sirm3d):

\[\sec^2x=1+\tan^2x=1+\frac{\sin^2x}{\cos^2x}=1+\frac{1}{\cos^2x} \sin^2x=1+\sec^2x \sin^2 x\]

OpenStudy (sirm3d):

left-to-right is not different from right-to-left.

OpenStudy (anonymous):

sec^2x=1/cos^2x=(cos^2x+sin^2x)/cos^2x=cos^2x/cos^2x+sin^2x/cos^2x=1+sin^2xsec^2x

OpenStudy (anonymous):

Oh! Thank you both so much! :)